Fermi National Laboratory
 
Operation Summary for February 1, 2012 - February 3, 2012
 

Wednesday February 1

The day shift began with day two of a seven-day shutdown of the accelerator complex to replace a Main Injector (MI) quadrupole magnet (Q403).

During the day shift, a Booster person reported that the East Gallery cable pull began at 8:35 AM. A Controls person began working on six front ends (FE) at 8:50 AM. Control personnel began replacing the Time Line Generator (TLG) hardware at 8:52 AM and completed the job at 9:22 AM; testing would be ongoing. At 10:10 AM, a Proton Source person turned off the 400 MeV area quadrupole magnet power supplies for 30 minutes. Operations reported at 11:58 AM that Linac was back on TCLK timing. Control personnel reported at 12:15 PM that the TLG was considered operational. Interlock personnel began conducting safety system tests at Meson enclosures M02, M03/4, and M05 at 1:19 PM. A Linac technician reported at 3:14 PM that he had completed his repair of a RF station (LRF1). A MI person reported at 3:25 PM that the MI work on the quadrupole magnet (Q403) replacement was progressing well; the welding was done and the vacuum pumpdown had started.

During the evening shift, LRF1 tripped off at 4:25 PM; it reset. A Linac person began a study at 5:30 PM.

Thursday January 2

The midnight shift began with day three of the seven-day accelerator complex shutdown.

During the midnight shift, at 12:23 AM, Operators found the Linac Klystron cooling water skid #1 was off; they turned on skid #2 in its place. At 5:13 AM, KRF1 tripped off; it reset. At 5:40 AM, KRF3 tripped off; it reset. At 7 AM, Neutron Therapy Facility (NTF) personnel began preparing to treat one patient later in the morning.

During the day shift, Operators reported at 8:15 AM that the scheduled two-hour MI-8 power outage had begun. Operators reported at 8:30 AM that the scheduled four-hour MI-12 power outage had begun. A Linac person began work on LRF5 at 10:35 AM. A Water Group person began work on the Booster glycol skid at 10:46 AM. A Linac person began work on the Buncher at 1:03 PM. Booster personnel began valving in magnets at 1:46 PM. FESS personnel reported at 1:47 PM that they were cleaning strainer baskets at Tevatron sectors E4, F1, F2, and F3; they had the job completed by 2:55 PM. Operators reported that the multiwires at MI-8 were off due to a bad power supply; this will be repaired on Friday. FESS personnel reported at 3:54 PM that they had repaired the air handler in the MI-60 North Power Supply room. Operators reported at the end of shift that the MI abort Lambertson bakeout had begun.

During the evening shift, FESS personnel reported at 4:36 PM that they completed cleaning the MI strainers. Operators completed a search and secure of the Booster at 6:45 PM. A Booster person began preparing for a hipot of Booster at 10:36 PM. Operators reported having all of the Booster power supplies on by 11:17 PM. Operations reported that the hipot was fine.

Friday February 3

The midnight shift began with day four of the seven-day accelerator complex shutdown.

During the midnight shift, at 12:02 AM, a Booster vacuum pump (B:IPRF22) tripped off; it reset. At 12:17 AM, LRF2 tripped off; it reset. At 1:55 AM, KRF1 tripped off; it reset. Operations prepared Booster for access at 6:16 PM. A Linac technician began working on LRF5 at 6:54 AM. At 7 AM, NTF personnel began preparing to treat two patients during the morning. Operations turned on the Muon Rings at 7:34 AM for Accumulator power supply evaluation studies.

Operations Coordinator (OpsCo) Plans for Today

The OpsCo reports that the Booster RF should be up later today and delivering beam to MiniBooNE by late tonight or early tomorrow midnight shift. MI might be ready to deliver beam by late on Tuesday, 2/7/12, or early on Wednesday, 2/8/12. The OpsCo agrees with plans by Muon Rings personnel to continue with power supply evaluations.
 

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 02/03/2012    email Fermilab
Security, Privacy, Legal  |  Use of CookiesFermi National Accelerator Laboratory